### v2

# An Improved Satterthwaite Effective Degrees of Freedom Correction for Weighted Syntheses of Variance

Abstract
This article presents an improved approximation for the effective degrees of freedom in the Satterthwaite (1941, 1946) method which estimates the distribution of a weighted combination of variance components The standard Satterthwaite approximation assumes a scaled chisquare distribution for the composite variance estimator but is known to be biased downward when component degrees of freedom are small. Building on recent work by von Davier (2025), we propose an adjusted estimator that corrects this bias by modifying both the numerator and denominator of the traditional formula. The new approximation incorporates a weighted average of component degrees of freedom and a scaling factor that ensures consistency as the number of components or their degrees of freedom increases. We demonstrate the utility of this adjustment in practical settings, including Rubin’s (1987) total variance estimation in multiple imputations, where weighted variance combinations are common. The proposed estimator generalizes and further improves von Davier’s (2025) unweighted case and more accurately approximates synthetic variance estimators with arbitrary weights.
Keywords: Satterthwaite approximation, effective degrees of freedom, variance components, weighted variances, multiple imputation, Rubin’s rules.
1 Introduction:
This article presents an improved approximation for the effective degrees of freedom in the Satterthwaite (1941, 1946) method, which is used for estimating the distribution of a weighted combination of variance components. The standard Satterthwaite approximation assumes a scaled chi-square distribution for the composite variance estimator but is known to be biased downward when component degrees of freedom are small. Building on recent work by von Davier (2025), we propose an adjusted estimator that corrects for this bias by modifying both the numerator and denominator of the traditional formula. The new approximation incorporates a weighted average of component degrees of freedom and a scaling factor that ensures consistency as the number of components or their degrees of freedom increase. We demonstrate the utility of this adjustment in practical settings, including Rubin’s (1987) total variance estimation in multiple imputation, where weighted variance combinations are common. The proposed estimator generalizes von Davier’s (2025) unweighted case and provides a more accurate approximation for synthetic variance estimators with arbitrary weights.
The Satterthwaite effective degrees of freedom approximation is defined for a synthesis of variance estimators. More specifically, for an expression
$\sigma_{*}^{2}=\sum_{k=1}^{K}w_{k}\sigma_{k}^{2}$
[…]
$\nu_{k}\frac{S_{k}^{2}}{\sigma_{k}^{2}}\sim\chi_{\nu_{k}}^{2}$
are used to calculate
$S_{*}^{2}=\sum_{k=1}^{K}S_{k}^{2}.$
The basic tenet of Satterthwaite (1941,1946) is that we may assume that
$\nu_{*}\frac{S_{*}^{2}}{\sigma_{*}^{2}}\sim\chi_{\nu_{*}}^{2}$
1.1 Things to Note
For $\chi^{2}-$distributed random variables it is well known that
$E\left[\nu_{*}\frac{S_{*}^{2}}{\sigma_{*}^{2}}\right]=\nu_{*}$
and
$Var\left[\nu_{*}\frac{S_{*}^{2}}{\sigma_{*}^{2}}\right]=2\nu_{*}.$
We will use some well known identities to arrive at the main result. First, recall that $Var\left(X\right)=E\left(X^{2}\right)-E\left(X\right)^{2}$ and that $Var\left(aX\right)=a^{2}Var\left(X\right)$ and that for independent $X_{1},...,X_{K}$we have $Var\left(\sum_{k}w_{k}X_{k}\right)=\sum_{k}w_{i}^{2}Var\left(X_{k}\right).$
The following equivalencies hold
$\nu_{*}^{2}\frac{Var\left(S_{*}^{2}\right)}{\left(\sigma_{*}^{2}\right)^{2}}=2\nu_{*}\leftrightarrow\nu_{*}=2\frac{\left(\sigma_{*}^{2}\right)^{2}}{Var\left(S_{*}^{2}\right)}$
and
$Var\left(S_{*}^{2}\right)=\sum_{k=1}^{K}w_{k}^{2}Var\left(S_{k}^{2}\right)=\sum_{k=1}^{K}w_{k}^{2}2\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}$
since
$Var\left(S_{k}^{2}\right)=2\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}$
which provides
$\nu_{*}=2\frac{\left(\sigma_{*}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}2\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\left(\sum_{k=1}^{K}w_{k}\sigma_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}}$
2 Satterthwaite Effective d.f. and Closer Approximations
Satterthwaite (1946) proposes
$\nu_{*}\approx\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{A}{B}\left(S_{\cdot}^{2},w_{\cdot}\right)$ (1)
as the effective degrees of freedom of $S_{*}^{2}.$
It was shown that this estuimator of the effective degrees of freedom for composite variances $S_{*}^{2}$ is biased downwards when the components $S_{k}^{2}$ have small or single degrees of freedom.
2.1 Some Properties of the d.f. Estimator
It is worth noting that the effect of the weights is invariant under rescaling and it can be assumed that all weights are postive, $w_{k}>0$, since zero weights would simply reduce the number of additive terms, and negative weights could lead to a negative variance. It can be shown easily that the same estimate will be obtained for two sets of weights $w_{\cdot},w_{\cdot}^{\#}$ with $w_{k}^{\#}=cw_{k}$ for all $k.$
$\frac{\left(\sum_{k=1}^{K}w_{k}^{\#}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}\left(w_{k}^{\#}\right)^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\left(\sum_{k=1}^{K}cw_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}c^{2}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{c^{2}\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{c^{2}\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}\right]}=\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}$
So we may define
$c=\min\left\{w_{k}^{\#}:k=1,\dots,K\right\}$
and can assume that we have at least one $w_{k}=1$, and $\forall k:w_{k}\geq 1$ from now on.
Trivially, the same property can be shown for the $S_{k}^{2}$, use $r=\min\left\{S_{k}^{2}:k=1,\dots,K\right\}$ and define
$R_{k}^{2}=\frac{1}{r}S_{k}^{2}$
so that there exists a $k^{{}^{\prime}}$with $R_{k^{{}^{\prime}}}^{2}=1$ and $R_{l}^{2}>R_{k^{{}^{\prime}}}^{2}$for all $l\neq k.$ Note that
$\frac{\left(\sum_{k=1}^{K}w_{k}R_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(R_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\left(\sum_{k=1}^{K}w_{k}\frac{1}{r}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(\frac{1}{r}S_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\frac{1}{r^{2}}\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\frac{1}{r^{2}}\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}\right]}=\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}.$
These results mean that the weights $w_{k}$ as well as the variance component estimators $S_{k}^{2}$ can be rescaled by multiplication with an arbitrary constant applied to all $k=1,\dots,K$ quantities.
3 Prior Improvements to the Satterthwaite Estimator
von Davier (2025) provided an improved approximation for small component d.f. $\nu_{k}$ and equal weights. The proposed effective d.f. is based on the observation that for small d.f. the simple replacement of $\sigma^{2}$ by substitution with $S^{2}$ does not hold in expectation. That is,
$E\left(\left[S_{k}^{2}\right]^{2}\right)\neq\left[\sigma_{k}^{2}\right]^{2}$
if $\nu_{k}<<\infty.$ For d.f. equals 1, it can be shown easily that $E\left[\left(S^{2}\right)^{2}\right]=\sigma^{4}E\left[Z^{4}\right]=3\sigma^{4}>\sigma^{4}.$
This leads to a simple adjustment for the denominator
$B_{new}=\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}$
so that for $\nu_{k}=1$ we adjust for $E\left(Z^{4}\right)=3$ while for $\nu_{k}\rightarrow\infty$ we have
$\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\rightarrow\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}.$
For the numerator $A,$ von Davier (2025) proposes the adjustment
$A_{new}=A\left(\frac{\left[K-1\right]\overline{\nu}+2}{\left[K-1\right]\overline{\nu}}\right)^{-1}=A\left(1+\frac{2}{\left[K-1\right]\overline{\nu}}\right)^{-1}=A\times f\left(\nu_{\cdot},K\right)$
with
$\overline{\nu}=\sum_{k=1}^{K}\frac{\nu_{k}}{K}$
being the arithmentic mean of the $\nu_{k}.$ The derivations by von Davier (2025) provide the following expression for the unweighted case:
$\nu_{*}\approx\frac{\left(\sum_{k=1}^{K}S_{k}^{2}\right)^{2}}{\left(1+\frac{K}{K-1}\frac{2}{\sum_{k}\nu_{k}}\right)\left[\sum_{k=1}^{K}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\right]}.$
4 An Improved Effective d.f. Estimator for the Weighted Case
Denoting the weighted average of the component d.f. $\nu_{k}$ by
$\overline{\nu_{w}}=\frac{\sum_{k}w_{k}\nu_{k}}{\sum w_{k}}$
it is proposed to utilize for cases where $K\geq 2$
$A_{new}=A\times\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)^{-1}.$
The above definition yield
$\nu_{*}\approx\frac{A_{new}}{B_{new}}=\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\right]}$ (2)
for a synthetic varince defined as weighted sums of variance components. Note that
$\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\right]}\rightarrow\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}\right]}$
as the (weighted) sum of the component degrees of freedom $\sum_{k}w_{k}\nu_{k}\rightarrow\infty$ grows. Similarly, with $K\rightarrow\infty$, we have that
$\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)\rightarrow 1$
even if $\nu_{k}=1$ remains for all $k.$
5 Further Improvement of the Approximation
The factor
$\left(1+\frac{2}{\left[K-1\right]\overline{\nu}}\right)^{-1}$
was derived based on the assumption that for $K=2$ and $\nu_{1}=\nu_{2}=1$ the maximum allowable approximate d.f. is $\nu_{*}=2.$ This value is obtained only if $S_{1}^{2}=S_{2}^{2}$, i.e., if the two are identical, or perfectly correlated.
However, a fundamental assumption is that the variance components $S_{k}^{2}$ are mutually independent (Satterthwaite, 1941; 1946), so that we will obtain $E\left(\nu_{*adj}\right)<K\nu=2=\nu_{1}+\nu_{2}$ in this case. A geometric argument (and simulation can also show applied researchers) shows that the expected value of $\nu_{*}$ for two independent variance components
$\sigma^{2}Z_{1}^{2},\sigma^{2}Z_{2}^{2}$
equals $E\left(\nu_{*adj}\right)=\sqrt{2}\approx 1.41<K\nu$ which is different from the limiting case obtained when assuming
$K,\nu\rightarrow\infty$
with $\nu_{k}=\nu$ for all $k=1,\dots,K$ where we find
$\lim_{K,\nu\rightarrow\infty}\nu\frac{\left(\sum_{k}S_{k}^{2}\right)^{2}}{\sum S_{k}^{4}}=\nu\frac{K^{2}\sigma^{4}}{K\sigma^{4}}=K\nu.$
It can be argued that it is desirable that $E\left(\nu_{*}\right)$ is not much smaller than $K=2$ in this case, while it should still converge to $K\nu$ in the case of equal d.f. for all components and with both $K,\nu$ growing.
The approach taken here is to argue that the integer part of the estimated d.f. $\nu_{*}$ would be used in most cases to provide a conservative estimate of the effective d.f. Then, an alternative upper limit $\nu_{max}=2.499\dots$ can be defined where we still have $int\left(\nu_{max}\right)=2$.
Then we obtain for $K=2,$ $\nu_{1}=\nu_{2}=1$ and $S_{1}^{2}=2_{2}^{2}$, the conditions that yield the maximum estimate,
$2.499\dots=\frac{3}{1+\frac{C}{K\nu}}\frac{\left(1^{2}+1^{2}\right)^{2}}{1^{4}+1^{4}}=\frac{6}{\left(K\nu+C\right)\frac{1}{K\nu}}=\frac{6}{\left(2+C\right)\frac{1}{2}}=\frac{12}{2+C}$
which leads to
$C=\frac{12}{2.499\dots}-2\approx 2.8.$
Note that this adjustment does not require the $K-1$ in the case where $C=2$. The proposed adjustment (von Davier), incidentally, can be replaced by
$A_{orig}=\left(1+\frac{2}{\left(K-1\right)\overline{\nu}}\right)^{-1}\approx\left(1+\frac{4}{K\overline{\nu}}\right)^{-1}$
it is proposed here to use
$A_{rev}=\left(1+\frac{2.8}{K\overline{\nu}}\right)^{-1}$
which yields
$\nu_{*}\approx\left(1+\frac{2.8}{K\overline{\nu_{w}}}\right)^{-1}\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}\frac{w_{k}^{2}S_{k}^{4}}{\nu_{k}+2}}$ (3)
6 Some Empirical Evidence for the Revised Adjustment
The following shows simulations using chisquare distributed variables (sums of squared standard normally distributed deviates). The columns define the degrees of freedom $\nu$ per component, the rows specify the number of components $K$. First, the original Satterthwaite (1941, 1946) formula is shown. It is known that this estimator of the effective degrees of freedom is biased when the component degrees of freedom are small. The results in table 2 and the following are based on 10000 replications each and the average in the table can be compared to the product of the row and column header value.
$K$ $\nu$ 1 2 3 4 5 10 20 40 80
2 1.42 3.15 4.98 6.85 8.76 18.49 38.32 78.18 158.08
3 1.81 4.23 6.90 9.63 12.43 26.99 56.51 116.25 236.17
4 2.19 5.29 8.74 12.38 16.05 35.33 74.77 154.42 314.18
5 2.56 6.37 10.61 15.10 19.76 43.67 92.99 192.68 392.21
10 4.37 11.56 19.83 28.53 37.74 85.50 183.88 383.09 782.55
20 7.84 21.67 37.93 55.41 73.44 168.89 365.79 764.04 1563.16
40 14.59 41.77 74.07 108.79 144.96 335.53 729.34 1525.83 3124.09
80 28.05 81.84 146.00 215.48 287.76 668.64 1456.81 3049.70 6246.00
160 54.72 162.04 289.96 428.71 573.51 1336.05 2911.30 6097.34 12489.83
Table 1: Averages of the estimated effective d.f. using the original Satterthwaite (1941, 1946) approach.
It can be seen that for component d.f. $\nu\in\left\{1,2,3,4,5\right\}$ the estimated effective d.f. are much smaller than the true value, that is, $\nu_{*,Satter}<<K\times\nu$. For example, the true value for $K=5$ and $\nu=3$ is $E\left(\nu_{*}\right)=15$ while the observed average is $M\left(\hat{\nu}_{*}\right)=10.61.$ Even for $\nu=10$ the estimator does not closely track the true value, for $K=4$ we find $M\left(\hat{\nu}_{*}\right)=35.33<40.$
The following table 2shows results for the new adjustment $1+\frac{2.8}{K\overline{\nu}}$ according to equation 3. As an example, for $\nu=10$ and $K=4$ we expect to see $E\left(\nu_{*}\right)=10\times 4=40$ while we observe $M\left(\hat{\nu}_{*}\right)=39.58\approx 40$, and for $\nu=3,K=5$ we obtain $M\left(\hat{\nu}_{*}\right)=14.88\approx 15$.
$K$ $\nu$ 1 2 3 4 5 10 20 40 80
2 1.77 3.69 5.63 7.60 9.57 19.49 39.35 79.29 159.29
3 2.80 5.80 8.76 11.71 14.70 29.58 59.38 119.26 239.28
4 3.85 7.85 11.87 15.81 19.72 39.58 79.49 159.31 319.24
5 4.90 9.97 14.88 19.96 24.74 49.63 99.46 199.27 399.30
10 10.17 20.22 30.10 40.03 50.13 99.79 199.65 399.38 799.22
20 20.53 40.45 60.41 80.20 99.97 199.84 399.60 799.60 1599.10
40 41.08 80.80 120.37 160.33 200.02 399.67 799.40 1599.11 3199.23
80 81.36 160.97 240.40 320.59 400.86 799.70 1599.37 3199.50 6398.96
160 161.44 320.64 481.18 640.53 800.88 1600.02 3199.55 6399.82 12799.47
Table 2: Averages of the estimated effective d.f. for the adjustment using $C=2.8$ and $K$
It can be seen that the observed average of the estimated degrees of freedom $M\left(\hat{\nu}_{*}\right)$ closely track the theoretical value $K\times\nu$ for most entries in this table.
The following table contains the results for the adjustment suggested by von Davier (2025), where $1+\frac{2}{\left(K-1\right)\overline{\nu}}$ was proposed. The differences are small, but it appears that the new adjustment is closer to the expected values by some small amount.
$K$ $\nu$ 1 2 3 4 5 10 20 40 80
2 1.41 3.15 4.96 6.84 8.76 18.50 38.29 78.20 158.07
3 2.72 5.66 8.59 11.53 14.50 29.39 59.25 119.19 239.07
4 3.97 7.99 11.99 15.91 19.89 39.73 79.54 159.50 319.38
5 5.14 10.19 15.14 20.14 25.07 49.96 99.69 199.52 399.62
10 10.66 20.81 30.71 40.61 50.48 100.22 200.09 400.03 799.87
20 21.16 41.14 60.98 80.97 100.84 200.47 400.25 800.16 1599.91
40 41.73 81.63 121.41 161.19 200.92 400.58 800.08 1600.15 3199.95
80 82.05 161.92 241.61 321.39 401.12 800.41 1600.35 3200.36 6399.96
160 162.39 322.04 481.80 641.26 800.84 1600.34 3199.96 6400.57 12799.94
Table 3: Averages of the estimated effective d.f. for the adjustment using $C=2$ and $K-1$
Especially for $\nu=1$, the cases shown in the first column, the adjustment provides a closer tracking of the expected $E\left(\nu_{*,new}\right)=K$, but also for varying $\nu$ and $K=2$ the new adjustment provides a slightly better approximation.
It could be argued that a rationale strategy could be to adjust for the case with $\nu=1$ and $K=2$ in an effort to match the observed value of
$E\left(\frac{\left[Z_{1}^{2}+Z_{2}^{2}\right]^{2}}{Z_{1}^{4}+Z_{2}^{4}}\right)\approx 1.4145$
in the table of average of estimated effective d.f.. This would require to find $C$ for which
$E\left(\nu_{*}\right)=2=\left(\frac{3}{1+\frac{C}{2}}\right)1.4145$
this leads to
$C=\frac{6-2\times 1.4145}{1.4145}\approx 2.2426.$
When using this adjustment $1+\frac{2.2426}{K\overline{\nu}}$, we obtain the following table that indeed shows the average estimated d.f. matches the expected value for the minimal case $K=2,\nu=1$ closely, we have $E\left(\nu_{*}\right)=M\left(\hat{\nu}_{K=2,\nu=1}\right)=2.00$. However, it appears that other entries in the table may show larger deviations compared to this proposed adjustment using $C=2.8.$
$K$ $\nu$ 1 2 3 4 5 10 20 40 80
2 2.00 4.02 6.03 8.03 10.04 19.96 39.89 79.84 159.80
3 3.11 6.16 9.16 12.15 15.12 30.05 59.97 119.85 239.85
4 4.22 8.27 12.31 16.27 20.27 40.15 79.95 159.88 319.82
5 5.32 10.36 15.36 20.38 25.33 50.10 99.94 199.87 399.87
10 10.68 20.79 30.57 40.57 50.49 100.28 200.05 399.91 799.90
20 21.20 40.89 60.79 80.91 100.66 200.28 400.13 799.98 1599.92
40 41.55 81.50 121.24 161.23 200.65 400.24 800.06 1599.97 3199.68
80 81.79 161.54 241.33 320.98 400.85 800.62 1599.77 3199.97 6399.89
160 161.62 321.70 481.37 641.00 800.72 1599.92 3200.27 6399.68 12799.68
Table 4: Averages of the estimated effective d.f. for the adjustment using $C=2.2426$ and $K$
How can we answer the question which of the thre adjustment is preferable overall? It appears that the adjustment using $C=2.2426$ more closely tracks the expected valkues for $K=2$ and varying $\nu\geq 1$, maybe not surprisingly, but at the same time deviations seem larger than the case using $C=2.8$ when looking at $K=3,4,\dots$ and small $\nu$. The overall quality of the approximations using different adjustments can be evaluated using a a pseudo chi-squared statistics as the table entries can be viewed as observed counts under an approximation while the expected counts are given by $K\times\nu.$ That is, a measure of average squared standardized deviation
$X_{C}^{2}=\sum_{\left\{K,\nu\right\}}\frac{\left(M_{C}\left(\hat{\nu}_{*}\right)-K\times\nu\right)^{2}}{K\times\nu}$
can be calculated over all entries in the table for each of the adjustment constants $C\in\left\{\left(2.8,K\right),\left(2,K-1\right),\left(2.2426,K\right)\right\}$. Note that larger deviations can typically be expected for small $K$ and small component $\nu$, and for large values of these the approximaytion is quite close, so that extending the table further beyond $K>160$ and $\nu>80$ would likely not change this pseudo $X^{2}$ measure much.
We obtain the following values for the three adjustnments compared in this paper based on the combinations of $K$and component $\nu$ shown in the tables above.
$C$ C $K_{adj}$ $X^{2}$
Satterthwaite (1941,1946) N/A N/A 859.8159
von Davier (2025) 2 K-1 1.7916
improved adjustment 2.8 2.8 K 0.3723
exact fit $(K=2,\nu=1)$ 2.2426 K 0.5376
Table 5: Pseudo $X^{2}$ values for the three adjustments under examination.
Table 5 shows that the lowest $X^{2}$ deviance among the three variants is obtained form $C=2.8$ and $K$ entering the adjustment without any correction. That is, with
$\overline{\nu}=\frac{\sum_{k=1}^{K}w_{k}\nu_{k}}{\sum_{k=1}^{K}w_{k}}$
it is proposed to use
$\nu_{*}\approx\left(1+\frac{2.8}{K\overline{\nu}}\right)^{-1}\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{S_{k}^{4}}{\nu_{k}+2}}$
as an estimator of effective degrees of freedom for the weighted case with varying numbers of components $K$ and for small and not-so-small component d.f. $\nu_{k}.$
7 Examples
7.1 Jackknifing and Balanced Repeated Replications (BRR)
In the case of single component d.f. $\nu_{k}=1$ and large(-ish) $K$ representing the (non-trivial) jackknife zones we have no weights to account for and obtain
$\nu_{*}\approx\frac{3}{1+\frac{2.8}{K}}\frac{\left(\sum_{k=1}^{K}\left[T_{\cdot}-T_{k}\right]^{2}\right)^{2}}{\left[T_{\cdot}-T_{k}\right]^{4}}$
since $\nu_{k}+2=3$ for all $k.$ Here, the $T_{k}$ are the pseudo-values calculated under the jackknife scheme of dropping units, and $T_{\cdot}=\sum_{k}T_{k}$. In the case of the BRR the $T_{k}$ are the balanced replicated estimates (of the half samples) obtained by adjusting the weights of the pairs and recalculating the statistics for each zone $K$.
7.2 Total Variance under Multiple Imputations
The total variance of an estimator when imputations is a prime example of a weighted variance estimate. The total variance as defined by Rubin (1987) is given by
$Var\left(total\right)=Var\left(sampling\right)+\frac{M+1}{M}Var\left(imputation\right)$
where the sampling variance may be estimated according to some resampling scheme (Johnson & Rust, 1992) and the imputation variance is the sample standard deviation across $M$ imputation based calculations of the same statistic.
In this case, the follwing weights are used:
$\left(w_{1},w_{2}\right)=\left(1,\frac{M+1}{M}\right)$
and equation 3 can directly be applied.
7.3 Welch Test
The modified Welch (1947) test uses the d.f. $\nu_{1}$ and $\nu_{2}$ of each of the variances of two samples to pool the variance. The Welch-Satterthwaite formula used in this case can be adjusted using the improved approximation proposed here. Then we have
$\nu_{*}\approx\left(1+\frac{2.8}{2\overline{\nu}}\right)^{-1}\frac{\left(\frac{1}{N_{1}}S_{1}^{2}+\frac{1}{N_{2}}S_{2}^{2}\right)^{2}}{\frac{S_{1}^{4}}{N_{1}^{2}\left(v_{1}+2\right)}+\frac{S_{1}^{4}}{N_{2}^{2}\left(v_{2}+2\right)}}$
where $S_{k}^{2}$ are the sample variances and weights
$w_{k}=\frac{1}{N_{k}}$
in this case, for the d.f. $\nu_{*}$ of the sample size weighted pooled variance used in the Welch test.
8 Conclusion
The above developments provide an improved Satterthwaite effective degrees of freedom approximation for the case of a synthetic variance estimator that utilizes weighted variance components. a prime example is the well known formula (Rubin, 1987) for the total variance for case that imputation are used to estimate sample statistics. The developments provided in this article generalize and further improve the adjusted Satterthwaite formula derived by von Davier (2025) for the unweighted case. It is proposed to estimate the effective degrees of freedom for the general case of $K\geq 2$ and $\nu_{k}\geq 1$ using equation 3