### v3

# An Improved Satterthwaite Effective Degrees of Freedom Correction for Weighted Syntheses of Variance

Abstract
This article presents an improved approximation for the effective degrees of freedom in the Satterthwaite (1941, 1946) method which estimates the distribution of a weighted combination of variance components The standard Satterthwaite approximation assumes a scaled chisquare distribution for the composite variance estimator but is known to be biased downward when component degrees of freedom are small. Building on recent work by von Davier (2025), we propose an adjusted estimator that corrects this bias by modifying both the numerator and denominator of the traditional formula. The new approximation incorporates a weighted average of component degrees of freedom and a scaling factor that ensures consistency as the number of components or their degrees of freedom increases. We demonstrate the utility of this adjustment in practical settings, including Rubin’s (1987) total variance estimation in multiple imputations, where weighted variance combinations are common. The proposed estimator generalizes and further improves von Davier’s (2025) unweighted case and more accurately approximates synthetic variance estimators with arbitrary weights.
Keywords: Satterthwaite approximation, effective degrees of freedom, variance components, weighted variances, multiple imputation, Rubin’s rules.
1 Introduction:
This article presents an improved approximation for the effective degrees of freedom in the Satterthwaite (1941, 1946) method, which is used for estimating the distribution of a weighted combination of variance components. The standard Satterthwaite approximation assumes a scaled chi-square distribution for the composite variance estimator but is known to be biased downward when component degrees of freedom are small. Building on recent work by von Davier (2025), we propose an adjusted estimator that corrects for this bias by modifying both the numerator and denominator of the traditional formula. The new approximation incorporates a weighted average of component degrees of freedom and a scaling factor that ensures consistency as the number of components or their degrees of freedom increase. We demonstrate the utility of this adjustment in practical settings, including Rubin’s (1987) total variance estimation in multiple imputation, where weighted variance combinations are common. The proposed estimator generalizes von Davier’s (2025) unweighted case and provides a more accurate approximation for synthetic variance estimators with arbitrary weights.
The Satterthwaite effective degrees of freedom approximation is defined for a synthesis of variance estimators. More specifically, for an expression
$\sigma_{*}^{2}=\sum_{k=1}^{K}w_{k}\sigma_{k}^{2}$
Satterthwaite (1941, 1946) argues that an approximate $\chi^{2}$ distribution can be assumed for variance estimators that are defined as a linear combination of variances, and proposes the following approximation of the degrees of freedom
$\nu_{*}\approx\frac{\left[\sum_{k}w_{k}S_{k}^{2}\right]^{2}}{\sum_{k}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}$
with components
$\nu_{k}\frac{S_{k}^{2}}{\sigma_{k}^{2}}\sim\chi_{\nu_{k}}^{2}$
that are used to calculate
$S_{*}^{2}=\sum_{k=1}^{K}S_{k}^{2}.$
The basic tenet of Satterthwaite (1941,1946) is that we may assume that
$\nu_{*}\frac{S_{*}^{2}}{\sigma_{*}^{2}}\sim\chi_{\nu_{*}}^{2}.$
1.1 Things to Note
For $\chi^{2}-$distributed random variables it is well known that
$E\left[\nu_{*}\frac{S_{*}^{2}}{\sigma_{*}^{2}}\right]=\nu_{*}$
and
$Var\left[\nu_{*}\frac{S_{*}^{2}}{\sigma_{*}^{2}}\right]=2\nu_{*}.$
We will use some well known identities to arrive at the main result. First, recall that $Var\left(X\right)=E\left(X^{2}\right)-E\left(X\right)^{2}$ and that $Var\left(aX\right)=a^{2}Var\left(X\right)$ and that for independent $X_{1},...,X_{K}$we have $Var\left(\sum_{k}w_{k}X_{k}\right)=\sum_{k}w_{i}^{2}Var\left(X_{k}\right).$
The following equivalencies hold
$\nu_{*}^{2}\frac{Var\left(S_{*}^{2}\right)}{\left(\sigma_{*}^{2}\right)^{2}}=2\nu_{*}\leftrightarrow\nu_{*}=2\frac{\left(\sigma_{*}^{2}\right)^{2}}{Var\left(S_{*}^{2}\right)}$
and
$Var\left(S_{*}^{2}\right)=\sum_{k=1}^{K}w_{k}^{2}Var\left(S_{k}^{2}\right)=\sum_{k=1}^{K}w_{k}^{2}2\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}$
since
$Var\left(S_{k}^{2}\right)=2\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}$
which provides
$\nu_{*}=2\frac{\left(\sigma_{*}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}2\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\left(\sum_{k=1}^{K}w_{k}\sigma_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(\sigma_{k}^{2}\right)^{2}}{\nu_{k}}}.$
2 Satterthwaite Effective d.f. and Closer Approximations
Satterthwaite’s (1946) estimator $\nu_{*,Satt}$ of the effective degrees of freedom for composite variances $S_{*}^{2}$ was found to be biased downwards when the components $S_{k}^{2}$ have small or single degrees of freedom $\nu_{k}<<\infty$ or even $\nu_{k}=1$ (e.g., Johnson & Rust, 1992; Qian, 1998).
2.1 Some Properties of the Weighted d.f. Estimator
It is worth noting that the effect of the weights is invariant under rescaling and it can be assumed that all weights are postive, $w_{k}>0$, since zero weights would simply reduce the number of additive terms, and negative weights could lead to a negative variance. It can be shown easily that the same estimate will be obtained for two sets of weights $w_{\cdot},w_{\cdot}^{\#}$ with $w_{k}^{\#}=cw_{k}$ for all $k.$
$\frac{\left(\sum_{k=1}^{K}w_{k}^{\#}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}\left(w_{k}^{\#}\right)^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\left(\sum_{k=1}^{K}cw_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}c^{2}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{c^{2}\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{c^{2}\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}\right]}=\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}$
So we may define
$c=\min\left\{w_{k}^{\#}:k=1,\dots,K\right\}$
and can assume that we have at least one $w_{k}=1$, and $\forall k:w_{k}\geq 1$ from now on.
The same property can be shown for the $S_{k}^{2}$, use $r=\min\left\{S_{k}^{2}:k=1,\dots,K\right\}$ and define
$R_{k}^{2}=\frac{1}{r}S_{k}^{2}$
so that there exists a $k^{{}^{\prime}}$with $R_{k^{{}^{\prime}}}^{2}=1$ and $R_{l}^{2}>R_{k^{{}^{\prime}}}^{2}$for all $l\neq k.$ Note that
$\frac{\left(\sum_{k=1}^{K}w_{k}R_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(R_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\left(\sum_{k=1}^{K}w_{k}\frac{1}{r}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(\frac{1}{r}S_{k}^{2}\right)^{2}}{\nu_{k}}}=\frac{\frac{1}{r^{2}}\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\frac{1}{r^{2}}\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}\right]}=\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}}.$
These results mean that the weights $w_{k}$ as well as the variance component estimators $S_{k}^{2}$ can be rescaled by multiplication with an arbitrary constant applied to all $k=1,\dots,K$ quantities.
3 Prior Improvements to the Satterthwaite Estimator
von Davier (2025) provided an improved approximation for small component d.f. $\nu_{k}$ and equal weights. The proposed effective d.f. is based on the observation that for small d.f. the simple replacement of $\sigma^{2}$ by substitution with $S^{2}$ does not hold in expectation. That is,
$E\left(\left[S_{k}^{2}\right]^{2}\right)\neq\left[\sigma_{k}^{2}\right]^{2}$
if $\nu_{k}<<\infty.$ For d.f. equals 1, it can be shown easily that $E\left[\left(S^{2}\right)^{2}\right]=\sigma^{4}E\left[Z^{4}\right]=3\sigma^{4}>\sigma^{4}.$
This leads to a simple adjustment for the denominator
$B_{new}=\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}$
so that for $\nu_{k}=1$ we adjust for $E\left(Z^{4}\right)=3$ while for $\nu_{k}\rightarrow\infty$ we have
$\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\rightarrow\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}.$
For the numerator $A,$ von Davier (2025) proposes the adjustment
$A_{new}=A\left(\frac{\left[K-1\right]\overline{\nu}+2}{\left[K-1\right]\overline{\nu}}\right)^{-1}=A\left(1+\frac{2}{\left[K-1\right]\overline{\nu}}\right)^{-1}=A\times f\left(\nu_{\cdot},K\right)$
with
$\overline{\nu}=\sum_{k=1}^{K}\frac{\nu_{k}}{K}$
being the arithmentic mean of the $\nu_{k}.$ The derivations by von Davier (2025) provide the following expression for the unweighted case:
$\nu_{*}\approx\frac{\left(\sum_{k=1}^{K}S_{k}^{2}\right)^{2}}{\left(1+\frac{K}{K-1}\frac{2}{\sum_{k}\nu_{k}}\right)\left[\sum_{k=1}^{K}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\right]}.$
4 An Improved Effective d.f. Estimator for the Weighted Case
Denoting the weighted average of the component d.f. $\nu_{k}$ by
$\overline{\nu_{w}}=\frac{\sum_{k}w_{k}\nu_{k}}{\sum w_{k}}$
it is proposed to utilize for cases where $K\geq 2$
$A_{new}=A\times\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)^{-1}.$
The above definition yield
$\nu_{*}\approx\frac{A_{new}}{B_{new}}=\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\right]}$ (1)
for a synthetic varince defined as weighted sums of variance components. Note that
$\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}+2}\right]}\rightarrow\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\left[\sum_{k=1}^{K}w_{k}^{2}\frac{\left(S_{k}^{2}\right)^{2}}{\nu_{k}}\right]}$
as the (weighted) sum of the component degrees of freedom $\sum_{k}w_{k}\nu_{k}\rightarrow\infty$ grows. Similarly, with $K\rightarrow\infty$, we have that
$\left(1+\frac{2}{(K-1)\overline{\nu_{w}}}\right)\rightarrow 1$
even if $\nu_{k}=1$ remains for all $k.$
5 Further Improvement of the Approximation
The adjustment factor
$\frac{3}{\left(1+\frac{C}{\left[K-p\right]\overline{\nu}}\right)}$
is a simple multiplicative term that is obtained if all component degrees of freedom are equal to one. Then we can write
$\nu_{*,2025}=\frac{3}{\left(1+\frac{C}{\left[K-p\right]\overline{\nu}}\right)}\nu_{*,Satt}$
with $p=1$ and assuming $\forall k:\nu_{k}=1=\nu$. This adjustment was derived based on the observation that $E\left(Z^{4}\right)=3$ and on the assumption that that for $K=2$ and $\nu_{1}=\nu_{2}=1$ the maximum allowable approximate d.f. is $\nu_{*}=2.$ This maximum value is obtained only if $S_{1}^{2}=S_{2}^{2}$, i.e., if the two are identical, or perfectly correlated. However, the regular case is that $S_{1}^{2},S_{2}^{2}$ are independent. For this minimal case, we find
$\frac{C_{1}}{\left[K-1\right]\overline{\nu}}=\frac{C_{2}}{K\overline{\nu}}$
for all $C_{2}=\frac{C_{1}K}{K-1}$ so we will focus on $p=0$ and adjustments of the type
$\hat{\nu}_{*,C}=\left(1+\frac{C}{K\overline{\nu}}\right)^{-1}\frac{\left[\sum_{k=1}^{K}w_{k}S_{k}^{2}\right]^{2}}{\sum_{k=1}^{K}\frac{S_{k}^{4}}{\nu_{k}+2}}$
in the following.
5.1 Matching the Expected Value
The case $K=2$ and $\nu_{1}=\nu_{2}=1$ will be retained as the finite value anchor for the adjustment as marks the smallest possible non-trivial combination of variance components. It can be shown that
$1\leq\frac{\left(S_{1}^{2}+S_{2}^{2}\right)^{2}}{S_{2}^{4}+S_{2}^{4}}\leq 2$
for any combination of $S_{1}^{2},S_{2}^{2}$. The true mean of the distribution must therefor be between $1$ and $2,$ so that we know the absolute difference is
$\left|E\left[\frac{\left(S_{1}^{2}+S_{2}^{2}\right)^{2}}{S_{2}^{4}+S_{2}^{4}}\right]-1.5\right|\leq 0.5.$
Picking the midpoint, $M=1.5$ for determining the correction constant $C,$we find
$2.252=\frac{3}{\left(1+\frac{C}{2}\right)}1.5\leftrightarrow 2\left(1+\frac{C}{2}\right)=4.5\leftrightarrow C=$ 2.25
as a value that should allow that approximated $M\left(\hat{\nu}_{*}\right)\approx 2.00$ for $K=2,\nu=1.$ That means, using these considerations, it is proposed to define
$\hat{\nu}_{*,2.25}=\frac{\left(\sum_{k}w_{k}S_{k}^{2}\right)^{2}}{\left(1+\frac{2.25}{K\overline{\nu}}\right)\left(\sum_{k}w_{k}^{2}\frac{\left[S_{k}^{2}\right]^{2}}{\nu_{k}+2}\right)}$
which retains the same limiting behavior for $K\rightarrow\infty$ and $\nu_{k}\rightarrow\infty$ we have
$\hat{\nu}_{*,2.25}\rightarrow\frac{\left(\sum_{k}w_{k}S_{k}^{2}\right)^{2}}{\left(\sum_{k}w_{k}^{2}\frac{\left[S_{k}^{2}\right]^{2}}{\nu_{k}}\right)}=\hat{\nu}_{*,Satt}.$
It could be argued that a rationale strategy could be to adjust for the case with $\nu=1$ and $K=2$ in an effort to match the observed average. A simulation of 4 million replications gives an observed value of
$E\left(\frac{\left[Z_{1}^{2}+Z_{2}^{2}\right]^{2}}{Z_{1}^{4}+Z_{2}^{4}}\right)\approx\frac{1}{4,000,000}\sum_{i=1}^{4,000,000}\frac{\left[Z_{1}^{2}+Z_{2}^{2}\right]^{2}}{Z_{1}^{4}+Z_{2}^{4}}=1.41425.$
This could be used to find a correctution term $C$ for which
$E\left(\nu_{*}\right)=2=\left(\frac{3}{1+\frac{C}{2}}\right)1.41425$
this leads to
$C=\frac{6-2\times 1.41425}{1.41425}\approx 2.2425.$
When using this adjustment $\left(1+\frac{2.2425}{K\overline{\nu}}\right)^{-1}$, we obtain the following table 4 that indeed shows the average estimated d.f. matches the expected value for the minimal case $K=2,\nu=1$ closely, we have $E\left(\nu_{*}\right)=M\left(\hat{\nu}_{K=2,\nu=1}\right)=2.00$.
However, two things to note here. First, this adjustment focuses on one case that does not seem to be very common, a case where there are only two components is not rare, of course, but cases where only two variance components are added together, and each component is essentially a squared normallly distributed variable, with one degree of freedom $\nu=1,$seem to be not those seen out in the wild often. Second, the behavior of the expression
$\frac{\left[Z_{1}^{2}+Z_{2}^{2}\right]^{2}}{Z_{1}^{4}+Z_{2}^{4}}=\hat{\nu}_{satt}$
for two independent normal deviates is shown in figure 1. It is obvious that the most frequently observed values are close to 1 and 2. This is the case because $Z_{1},Z_{2}$ are independent and in the more likely case that $Z_{i}<<Z_{j}$ then $\hat{\nu}_{satt}\approx 1$ while it is less likely that $Z_{1}\approx Z_{2}$ which leads to $\hat{\nu}_{satt}\approx 2$. So it would seem that focusing on a value that matches the empirical $1.41425$ for a single $K,\nu$ combinations is not the most useful strategy.
Figure 1: Empirical density of $R$ based on 4,000,000 replicates of two independent normal deviates.
This lead to a consideration that goes beyond finding an adjustment that exactly matches the theoretical value for one of the potential combinations of $K,\nu$. Instead, the next section explores how a constant can be found that minimizes the difference from the expected result $K\nu$ for a range of components and component degrees of freedom.
5.2 Minimizing Squared Residuals
In this section, we will step away from the somewhat reductionist view that limits the problem to only matching the expected value $E\left(\nu_{*}\right)=K\nu$ for $K=2$ and equal component degrees of freedom. Instead, this section looks at the small d.f. behavior of a range of adjustment factors $C\in\Omega_{C}=\left]2.0,3.2\right[$(why we do not look outside that range will be clear when examining figure ). We aim to cover the small d.f. and low component count cases where the adjustment to the effective d.f. are expected to have the largest effects. The simulations use the best case scenario where we know each component is $\chi^{2}$ distributed with d.f. $\nu_{k}$, that is, each variance component can be written as
$S_{k}^{2}=\sum_{j=1}^{\nu_{k}}Z_{j}^{2}$
with $Z_{j}\sim N\left(0,1\right)$, so that $E\left(S_{k}^{2}\right)=\nu_{k}$ and $Var\left(S_{k}^{2}\right)=2\nu_{k}$. In this case, for $S_{*}^{2}=\sum_{k=1}^{K}S_{k}^{2}$ we can determine the expected value is $E\left(\nu_{*}\right)=K\times\nu$. Within the range of potential constants $C\in\Omega_{C}$, the choice will be made based on an optimality criterion
$X_{C}^{2}=\sum_{K\in\Omega_{K}}\sum_{\nu\in\Omega_{\nu}}\frac{\left[M\left(\hat{\nu}_{*,C}\right)-K\nu\right]^{2}}{K\nu}$
where the range of components $K$ and component degrees of freedom $\nu$ is defined as
$K\in\Omega_{K}=\left\{2,3,\dots,K_{max}\right\}\,\,\mathrm{and}\,\,\nu\in\Omega_{\nu}=\left\{1,2,\dots,\nu_{max}\right\}.$
The quantity $X_{C}^{2}$ can be understood as a pseudo $\chi^{2}$ measure of deviation, which is minimized by finding
$C_{opt}=\arg\min_{C\in\Omega_{C}}\left\{X_{C}^{2}\right\}$
to determine an optimal choice for this range of components $K$ and compoment d.f. $\nu$. In order to obtain estimates of the expected effective d.f. $E\left(\nu_{*,C}\right)$ a number of $r=10000$ replicates was used to calculate
$\frac{1}{10000}\sum_{i=1}^{10000}\hat{\nu}_{*,C}=M\left(\hat{\nu}_{*,C}\right)\approx E\left(\nu_{*,C}\right).$
In terms of the practical implementation of the search, the optimal constant $C_{opt}$ was determined using a line search for different values of $\left(K_{max},\nu_{max}\right)$ and it was monitored if the choice converges for these different upper limits. More specifically, the following combinations were checked:
$\left(K_{max},\nu_{max}\right)\in\left\{\left(5,5\right),\left(10,10\right),\left(20,20\right),\left(30,30\right),\left(40,40\right),\left(50,50\right)\right\}$
Then, $C_{opt}$ is used to define the improved effective degree of freedom as
$\nu_{*}\approx\left(1+\frac{C_{opt}}{K\overline{\nu_{w}}}\right)^{-1}\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}\frac{w_{k}^{2}S_{k}^{4}}{\nu_{k}+2}}.$ (2)
Note that, as before, the correction factor $\left(1+\frac{C_{opt}}{K\overline{\nu}}\right)^{-1}$is only important for small $K,\overline{\nu}$. The limiting case is obtained when assuming
$K,\overline{\nu}\rightarrow\infty$
with $\nu_{k}=\nu$ for all $k=1,\dots,K$. In that case we find the well known result
$\lim_{K,\nu\rightarrow\infty}\left(1+\frac{C}{K\nu}\right)^{-1}\left[\nu+2\right]\frac{\left(\sum_{k}S_{k}^{2}\right)^{2}}{\sum S_{k}^{4}}=\nu\frac{K^{2}\sigma^{4}}{K\sigma^{4}}=K\nu.$
That is why it is sufficient to examine a set of limited ranges $\Omega_{K}\times\Omega_{\nu}=\left\{2,3,\dots,K_{max}\right\}\times\left\{1,2,\dots,\nu_{max}\right\}$ as with growing $K,\nu$ the adjustment will converge to the original Satterthwaite (1946) equation.
6 Empirical Evidence for the Revised Adjustment
This section explains how the value $C_{opt}$ was chosen to provide an optimal adjustment for a range of small to medium values for $K$ and $\nu_{k}$ and how the improvement of the Satterthwaite (1941, 1946) effective degrees of freedom fares compared to the original, and in comparison to the improvement proposed by von Davier (2025) as well as the contender that aims to match the expected degrees of freedom using $C=2.25,$to the theoretical value for $K=2$ and $\nu_{1}=\nu_{2}$=1.
6.1 Determining $C_{opt}$
Figure 2 shows the graph for $X^{2}$ as a function of $C$ for the different selections of $K_{max},\nu_{max}.$ It appears that the minimum value starts out at around $C_{opt}\left(5,5\right)\approx 2.43$ and converges as $K_{max}$ and $\nu_{max}$ increase to about $2.66\leq C_{opt}\leq 2.68$.
Figure 2: Observed $X^{2}$ discrepancies for different values of $K_{max},\nu_{max}$
The way the optimal value was determined as $C_{opt}\approx 2.67$ can be explained based on table 1. The growth of the values of $C_{opt}$ as $K_{max},\nu_{max}$ grow is negatively accelerated and for the difference between $\left(40,40\right)$ and $\left(50,50\right)$ we see that $C_{opt}$only increases by 0.29% from $2.6511$ to $2.6588$The percentage increase seems to be cut in half for each increase of $K_{max},\nu_{max}$ by 10. Extrapolating this pattern to 160 shows that the expected growth of $C_{opt}$ is limited by $2.67$.
$K,\nu$ $C_{opt,K,\nu}$ $C_{opt}\frac{next}{previous}$ Change
Observed Values
5,5 2.4340 NA NA
10,10 2.5391 1.043171 4.3171
20,20 2.6094 1.027674 2.7674
30,30 2.6367 1.010460 1.0460
40,40 2.6511 1.005498 0.5498
50,50 2.6588 1.002896 0.2896
Predicted Values
60,60 2.6628 1.001500 0.1500
70,70 2.6650 1.000777 0.0777
80,80 2.6661 1.000402 0.0402
90,90 2.6667 1.000208 0.0208
100,100 2.6671 1.000108 0.0108
110,110 2.6673 1.000056 0.0056
120,120 2.6674 1.000029 0.0029
130,130 2.6674 1.000015 0.0015
140,140 2.6674 1.000008 0.0008
150,150 2.6674 1.000004 0.0004
160,160 2.6675 1.000002 0.0002
Table 1: Optimal values and growth rates for $C_{opt}$ by different combinations of $K_{max},\nu_{max}.$
6.2 Comparing the Original with Old and New Adjustments
The following tables show the averages for the unadjusted $\hat{\nu}_{*,Satter}$, for the von Davier (2025) adjustment $\hat{\nu}_{*,vD}$, and for the proposed improved adjustment using $C_{opt}=2.67,$ as well as for $\hat{\nu}_{*,2.25}$ to match the $K=2,\nu=1$ case exactly. The following tables show results for the average estimates of $\hat{\nu}_{*}$over 10000 replications and for the following subset of component counts and component degrees of freedom
$K\in\Omega_{K/2}=\left\{2,4,6,8,10,20,40,160\right\}$
and
$\nu\in\Omega_{\nu/2}=\left\{1,3,5,7,9,15,30,80\right\}$
This selection was chosen make it possible to show tables on a single page. Their associated $X^{2}$ sums of standardized squared deviations for each case are given in an additional table based on the full set of $K,\nu$ combinations. The results in tables2,3,4 and 5 are based on 10000 replications each and the average in the table can be compared to the product of the row and column header value.
In the tables, the average of the estimated d.f. are shown. The columns specify the degrees of freedom $\nu$ per component, the rows specify the number of components $K$. First, the original Satterthwaite (1941, 1946) formula is shown. It is known that this estimator of the effective degrees of freedom is biased when the component degrees of freedom are small.
$K\backslash\nu$ 1 3 5 7 9 15 30 80
2 1.42 4.98 8.77 12.63 16.53 28.40 58.22 158.10
4 2.19 8.79 16.06 23.67 31.37 54.98 114.50 314.24
6 2.93 12.45 23.36 34.67 46.20 81.49 170.80 470.27
8 3.66 16.14 30.54 45.51 60.94 107.98 227.04 626.48
10 4.36 19.86 37.77 56.43 75.70 134.34 283.21 782.49
20 7.81 37.96 73.39 110.88 149.31 266.92 564.55 1563.00
40 14.63 74.08 145.17 219.90 296.66 531.68 1127.06 3124.05
160 54.79 290.21 573.48 873.04 1180.50 2119.89 4502.64 12489.85
Table 2: Averages of the estimated effective d.f. using the original Satterthwaite (1941, 1946) approach.
For the original Satterthwaite (1941, 1946) approach, the empirical averages are given in table 2. It can be seen that for component d.f. $\nu\in\left\{1,3,5,7,9\right\}$ the estimated effective d.f. $\hat{\nu}_{*}$are much smaller than the true value, that is, $\hat{\nu}_{*,Satter}<<K\times\nu$. For example, the true value for $K=6$ and $\nu=3$ is $E\left(\nu_{*}\right)=18$ while the observed average is $M\left(\hat{\nu}_{*}\right)=12.45.$ Even for $\nu=9$ the estimator does not closely track the true value, for $K=4$ we find $M\left(\hat{\nu}_{*}\right)=31.37<36.$ Maybe surprisingly, even for $\nu=30,K=160$ we obtain $M\left(\hat{\nu}_{*}\right)=4502.64<4800=30\times 160.$
The following table 3 contains the results for the adjustment suggested by von Davier (2025), where $\left(1+\frac{2}{\left(K-1\right)\overline{\nu}}\right)^{-1}$ was proposed. The differences are small, but it appears that the new adjustment using $C_{opt}$ is closer to the expected values by some amount.
$K\backslash\nu$ 1 3 5 7 9 15 30 80
2 1.42 4.97 8.77 12.61 16.55 28.36 58.20 158.04
4 3.95 11.95 19.88 27.76 35.71 59.68 119.49 319.42
6 6.28 18.35 30.15 42.11 54.06 89.99 179.74 479.62
8 8.48 24.58 40.46 56.34 72.19 120.17 239.90 639.62
10 10.71 30.75 50.51 70.34 90.30 150.06 300.02 799.78
20 21.23 61.00 100.87 140.91 180.51 300.37 600.03 1599.88
40 41.71 121.32 200.89 280.83 360.65 600.35 1200.25 3199.69
160 162.21 481.46 801.11 1120.34 1440.82 2400.28 4800.55 12800.27
Table 3: Averages of the estimated effective d.f. for the von Davier (2025) adjustment using $C=2$ and $K-1$
It can be seen that the expected value for $K=2$ and $\nu=1$ is the same as for the original Satterthwaite equation, but otherwise the adjustment proposed by von Davier (2025) seems to track the theoretical values of $K\times\nu$ closely. However, the next two tables will show that closer approximations are possible.
The next table shows the adjustment for $C=2.25$ which attempts to match $K\nu=2\times 1$ closely for the limiting case with only two components and single d.f. It can be seen that the expected value for this adjusted Satterthwaite formula is indeed $2.00$ as given in table 4.
$K\backslash\nu$ 1 3 5 7 9 15 30 80
2 2.00 6.05 10.01 13.99 18.00 29.93 59.83 159.83
4 4.22 12.28 20.23 28.21 36.09 60.06 119.97 319.78
6 6.41 18.53 30.32 42.21 54.20 90.03 179.89 479.87
8 8.51 24.59 40.46 56.36 72.22 120.09 239.94 639.94
10 10.72 30.62 50.56 70.44 90.31 150.17 299.88 799.80
20 21.11 60.89 100.71 140.50 180.28 300.02 600.16 1600.04
40 41.63 121.18 200.69 280.73 360.63 600.30 1199.88 3199.90
160 161.96 481.45 801.08 1120.87 1440.76 2400.07 4800.32 12799.96
Table 4: Averages of the estimated effective d.f. for the adjustment using $C=2.25$ and $K$
Table 5shows results for the proposed improved adjustment using the factor $\left(1+\frac{2.67}{K\overline{\nu}}\right)^{-1}$ according to equation 2. As an example, for $\nu=9$ and $K=6$ we expect to see $E\left(\nu_{*}\right)=9\times 6=54$, and indeed we observe $M\left(\hat{\nu}_{*}\right)=53.76\approx 54$, and for $\nu=3,K=8$ we obtain $M\left(\hat{\nu}_{*}\right)=24.11\approx 24$.
$K\backslash\nu$ 1 3 5 7 9 15 30 80
2 1.81 5.72 9.69 13.62 17.58 29.55 59.46 159.40
4 3.94 11.94 19.90 27.78 35.73 59.68 119.48 319.35
6 6.07 18.11 29.96 41.95 53.76 89.59 179.55 479.28
8 8.25 24.11 40.11 56.01 71.86 119.70 239.57 639.33
10 10.37 30.39 50.12 70.01 89.84 149.71 299.43 799.48
20 20.78 60.49 100.06 139.91 180.18 299.46 599.77 1599.50
40 41.13 120.57 200.41 280.18 359.87 599.68 1199.48 3199.39
160 161.98 481.17 800.72 1119.78 1440.02 2399.65 4799.92 12799.01
Table 5: Averages of the estimated effective d.f. for the adjustment using $C=2.67$ and $K$
It can be seen that the observed average of the estimated degrees of freedom $M\left(\hat{\nu}_{*}\right)$ closely track the theoretical value $K\times\nu$ for most entries in this table. Especially for $\nu=1$, the cases shown in the first column, the new adjustment provides a closer tracking of the expected $E\left(\nu_{*,new}\right)=K$, but also for varying $\nu$ and $K=2$ the new adjustment provides a slightly better approximation than the one using $C=2.25$.
The overall quality of the approximations using different adjustments can be evaluated using the pseudo $\chi^{2}$ statistics $X^{2}$ calculated for each of the adjustment constants $C\in\left\{\left(2.67,K\right),\left(2,K-1\right),\left(2.25,K\right)\right\}$ as well as the original Satterthwaite (1941, 1946) formula. Note that larger deviations can typically be expected for small $K$ and small component $\nu$, and for large values of these the approximaytion is quite close, so that extending the table further beyond $K>160$ and $\nu>80$ would likely not add much to the pseudo $X^{2}$ measure. As expected, the table 6shows that proposed adjustment is indeed the one that shows the smallest discrepancy among the four.
$C$ $K_{adj}$ $X^{2}$
Satterthwaite (1941,1946) N/A N/A 13.27251
von Davier (2025) 2 K-1 0.31631
Assume $E\left(\hat{\nu}_{*}\right)\approx 1.5$ for $(K=2,\nu=1)$ 2.25 K 0.06412
Optimal $\left(K_{max},\nu_{max}\right)$ Adjustment $C_{opt}$ 2.67 K 0.03632
Table 6: Pseudo $X^{2}$ values for the original Satterthwaite estimator, and the three adjustments under examination.
We can confirm that the for the three adjustnments compared in this paper based on the combinations of $K$ and component $\nu$ shown in the tables above that the adjustment using $C_{opt}$ leads to the overall smallest $X^{2}$. That is, with
$\overline{\nu}=\frac{\sum_{k=1}^{K}w_{k}\nu_{k}}{\sum_{k=1}^{K}w_{k}}$
it is proposed to use
$\nu_{*}\approx\left(1+\frac{2.67}{K\overline{\nu}}\right)^{-1}\frac{\left(\sum_{k=1}^{K}w_{k}S_{k}^{2}\right)^{2}}{\sum_{k=1}^{K}w_{k}^{2}\frac{S_{k}^{4}}{\nu_{k}+2}}$ (3)
as an estimator of effective degrees of freedom for weighted and unweighted cases with varying numbers of components $K$ and for small and not-so-small component d.f. $\nu_{k}.$
7 Examples
7.1 Jackknifing, Balanced Repeated Replications (BRR)
In the case of single component d.f. $\nu_{k}=1$ and large(-ish) $K$ representing the (non-trivial) jackknife zones we have no weights to account for and obtain
$\nu_{*}\approx\frac{3}{1+\frac{2.8}{K}}\frac{\left(\sum_{k=1}^{K}\left[T_{\cdot}-T_{k}\right]^{2}\right)^{2}}{\left[T_{\cdot}-T_{k}\right]^{4}}$
since $\nu_{k}+2=3$ for all $k.$ Here, the $T_{k}$ are the pseudo-values calculated under the jackknife scheme of dropping units, and $T_{\cdot}=\sum_{k}T_{k}$. In the case of the BRR the $T_{k}$ are the balanced replicated estimates (of the half samples) obtained by adjusting the weights of the pairs and recalculating the statistics for each zone $K$.
7.2 Total Variance under Multiple Imputations
The total variance of an estimator when imputations is a prime example of a weighted variance estimate. The total variance as defined by Rubin (1987) is given by
$Var\left(total\right)=Var\left(sampling\right)+\frac{M+1}{M}Var\left(imputation\right)$
where the sampling variance may be estimated according to some resampling scheme (Johnson & Rust, 1992) and the imputation variance is the sample standard deviation across $M$ imputation based calculations of the same statistic.
In this case, the follwing weights are used:
$\left(w_{1},w_{2}\right)=\left(1,\frac{M+1}{M}\right)$
and equation 2 can directly be applied.
7.3 Welch (1947) Test
The modified Welch (1947) test uses the d.f. $\nu_{1}$ and $\nu_{2}$ of each of the variances of two samples to pool the variance. The Welch-Satterthwaite formula used in this case can be adjusted using the improved approximation proposed here. Then we have
$\nu_{*}\approx\left(1+\frac{2.8}{2\overline{\nu}}\right)^{-1}\frac{\left(\frac{1}{N_{1}}S_{1}^{2}+\frac{1}{N_{2}}S_{2}^{2}\right)^{2}}{\frac{S_{1}^{4}}{N_{1}^{2}\left(v_{1}+2\right)}+\frac{S_{1}^{4}}{N_{2}^{2}\left(v_{2}+2\right)}}$
where $S_{k}^{2}$ are the sample variances and weights
$w_{k}=\frac{1}{N_{k}}$
in this case, for the d.f. $\nu_{*}$ of the sample size weighted pooled variance used in the Welch test.
8 Conclusion
The above developments provide an improved Satterthwaite effective degrees of freedom approximation for the case of a synthetic variance estimator that utilizes weighted variance components. a prime example is the well known formula (Rubin, 1987) for the total variance for case that imputation are used to estimate sample statistics. The developments provided in this article generalize and further improve the adjusted Satterthwaite formula derived by von Davier (2025) for the unweighted case. It is proposed to estimate the effective degrees of freedom for the general case of $K\geq 2$ and $\nu_{k}\geq 1$ using equation 2